OpenStudy (anonymous):

A triangle has an area of 24 square cm and a base of 4 cm. Which is the height of the triangle?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

24/4=6

OpenStudy (anonymous):

\[A = \frac{ 1 }{ 2 } \times b \times h\]

Directrix (directrix):

1/2*6*4 does not equal 24

OpenStudy (anonymous):

yeah i know @Directrix that's why i put the area of a triangle formula.

OpenStudy (anonymous):

@hahd you are wrong bro.

OpenStudy (anonymous):

\[24 = \frac{ 1 }{ 2 } \times 4 \times h\]

OpenStudy (anonymous):

woops lol

OpenStudy (anonymous):

my credibility just went down the drain

OpenStudy (anonymous):

i just woke up so pls

OpenStudy (anonymous):

\[24 = 2 \times h\]

OpenStudy (anonymous):

\[12 = h\]